OpenStudy (anonymous):

Select one of the factors of x3y2 + 8xy2 - 5x2 - 40. Can someone help me figure this out?

OpenStudy (anonymous):

answer choices (xy2 + 5) (x2 + 4) (xy2 - 5) (x2 - 8)

OpenStudy (anonymous):

Try solving by grouping! Like so, x3y2+8xy2-5x2-40 x3y2-5x2+8xy2-40 (rearranged so degree of x is in descending order) (x3y2-5x2)+(8xy2-40) (Grouping) x2(xy2-5)+8(xy2-5) (taking out factors) (x2+8)(xy2-5) (factored) Check and see if any of those factors match!

OpenStudy (anonymous):

i'm still kinda confused on why you rearranged it @NCount

OpenStudy (neer2890):

\[x ^{2}(xy ^{2}-5)+8(xy ^{2}-5)=(x ^{2}+8)(xy ^{2}-5)\]

OpenStudy (anonymous):

Select one of the factors of 3x2 + 4x - 4. (x + 4) (3x + 2) (3x - 2) (x - 2) @neer2890

OpenStudy (anonymous):

i'm just having trouble one these

OpenStudy (anonymous):

I rearranged it in an order so I could group certain expressions together. Notice how when I pulled out the GCF of both expressions, they yielded the same expression.

OpenStudy (neer2890):

\[3x ^{2}+6x-2x-4=0\] 3x(x+2)-2(x+2)=0 (3x-2)(x+2)=0

OpenStudy (anonymous):

okay thank you can you help with the other one too? @NCount

OpenStudy (anonymous):

Sure thing, yet it seems like neer 2890 beat me too it. :P What he did there is realized that the 4x could be separated into the 6x and −2x (since when you add them together, they yield 4x) and used the same principle.

OpenStudy (anonymous):

thats just confusing to me but i wanna know how to do it instead of just getting the answers...

OpenStudy (anonymous):

It definitely takes practice in learning to recognize when to manipulate the problem as such.

OpenStudy (anonymous):

yea i know i'm just confused

OpenStudy (anonymous):

i still dont get where the 6x and -2x came from @NCount

OpenStudy (neer2890):

ok. so you can do this by two methods 1. To find roots of quadratic equation\[ax ^{2}+bx+c=0\] \[x=\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (neer2890):

2.in \[ax ^{2}+bx+c=0\] mutiply ax^2 and c, the value you got then you have to factor that value into two parts such that their addition becomes bx and their multiplication becomes cax^2., for e.g.,\[3x ^{2}+4x-4=0\] mutiply 3x^2 and -4= -4*3x^2=-12x^2 now factor -12x^2 into two parts such that it becomes +4x i chose 6x and -2x because 6x-2x=4x and 6x*(-2x)=-12x^2